# Why In-Context Learning Transformers are Tabular Data Classifiers

### v1

@@ 5.1 Introduction of the Benchmark Datasets
TabPFN (original) Zero-shot 0.530 0.587
TabPFN (original) Fine-tune 0.716 0.738
TabPFN (retrained) Zero-shot 0.207 0.351
TabPFN (retrained) Fine-tune 0.731 0.750
TabForest Zero-shot 0.208 0.350
TabForest Fine-tune 0.829 0.804
TabForestPFN Zero-shot 0.402 0.570
TabForestPFN Fine-tune 0.848 0.835
The TabZilla [32] benchmark tests 20 algorithms on 176 classification datasets with sizes ranging from 32 observations to over a million. We selected 94 out of 176 datasets, see appendix A.5. The medium size benchmark which we refer to as WhyTrees [15] consists of 23 classification datasets with 2923 to a maximum of 10,000 observations. The benchmark is split into 7 datasets with only numerical features and 16 datasets with both numerical and categorical features.
5.2 Main Results of TabForestPFN
The results on the TabZilla benchmark are shown in Table 3, see Appendix A.6 for alternative presentations. For the WhyTrees benchmark, the comparison of fine-tuned TabForestPFN with the benchmark algorithm is shown in Figure 4, and the comparison with other ICL-transformer variants in Table 2.
Table 3: Main Results on TabZilla. N. Accuracy stands for Normalized accuracy. Rank compares the relative rank of a method compared to all other methods on that dataset. Only the 5 best methods out of 26 are shown. Full table with all methods is presented in Appendix A.6
@@ 5.2 Main Results of TabForestPFN
TabForestPFN - Fine-tune 1 24 8.0 6.0 0.680 0.663
TabPFN (retrained) - Fine-tune 1 26 8.3 7.8 0.678 0.700
CatBoost 1 22 8.9 7.8 0.676 0.663
XGBoost 1 23 9.2 8.2 0.674 0.671
TabPFN (original) - Fine-tune 1 25 9.3 8.0 0.670 0.677
@@
In these figures and tables, we see that the fine-tuned TabForestPFN outperforms all other methods on TabZilla and matches the performance of XGBoost with a thousand runs on mixed features and a hundred runs on numerical features on WhyTrees. For both benchmarks, we can also see that there is a significant gap between the zero-shot performance of our retrained TabPFN and the original TabPFN. We attribute this phenomenon to the training settings, see Appendix A.4
5.3 Similarity of TabPFN and TabForest
Now that we have established the strong performance of TabForestPFN, we take a closer look at the behavior of TabPFN versus TabForest. In Tables 4 and 6, we see that there is no clear favorite between the fine-tuned versions of TabPFN and TabForest: TabForest is the best on WhyTrees, while TabPFN is favored on TabZilla. Furthermore, TabForestPFN only has a marginal improvement over TabPFN on TabZilla, and similarly only a small performance increase over TabForest on WhyTrees.
In Figure 5 we show the performance of TabPFN, TabForest and TabForestPFN on individual datasets from the WhyTrees benchmark. Remarkably, all three methods often show similar performance across datasets, even though TabPFN and TabForest are pretrained on completely different synthetic data. These results suggest the TabPFN’s performance does not come from the fact that the synthetic dataset is a "prior". Instead the results suggest the performance is a property of an ICL-transformer’s learning process.
Furthermore, we see that there are some differences on some individual datasets, such as the eye movements and, to a lesser extent, the bank-marketing and MiniBooNE datasets. Although we cannot explain why these differences exist, it does prompt us to ensemble the synthetic datasets to create TabForestPFN, which seems to mimic the best method of the two.
Figure 5: Comparison of TabForest, TabPFN and TabForestPFN on 10 datasets of the WhyTrees benchmark. Performance of different ICL-transformers is often similar. All the benchmark algorithms are greyed out. See Appendix A.7 for all WhyTrees datasets and for colored benchmark lines.
@@ 5.4 Complexity of ICL-Transformers’ Decision Boundaries
In the previous section, we have seen TabForest and TabPFN show similar performance behavior. Now we take a look at their decision boundaries. Repeating the analysis of the WhyTrees’ authors [15], we use Electricity dataset (OpenML ID: 44156) to predict a target on two features.
To capture the complexity of the decision boundary, we define the complexity score $V$. We split the feature space into a total of $n$ grid cells where each cell has a predicted probability of class one $p_{ij}$ for grid cell indices $(i,j)$. The complexity score $V$ is defined as the sum of absolute values between neighbor cells:
@@
We plot the results in Figure 1. We see that when fine-tuning, both TabPFN and TabForest can create decision boundaries that are more complex than regular MLPs. The complexity of the decision boundaries was one of the characteristics that explained why tree-based methods outperformed neural networks [15]. These results suggest ICL-transformers can also create complexity in the decision boundaries. In Appendix A.10 we also show that the complexity scales with the dimension of the ICL-transformer.
In our intuition, the ICL-transformer learns how to create these decision boundaries during pretraining. Different parameters of the network might be responsible for different parts of the decision boundary. During zero-shot, the ICL-transformer can then create the decision boundary out of these parts. During fine-tuning, the effect is stronger, as different parts come to the surface while irrelevant parts are tuned down.
@@ 5.5 Ablation of the Forest Dataset Generator
In Section 4.1 we discussed two ways to influence the complexity of the forest dataset generator: the tree depth and the base size, which is the number of observations to fit the tree algorithm. We expect the performance of TabPFN to increase when the complexity of the forest dataset generator increases.
In Figure 6 we show the results of pretraining different settings of base size and maximum tree depth on the WhyTrees benchmark. The tree depth is set to 1-25 as the base size changes, and the base size is fixed to 1024 as the tree depth changes. When scaling up the base size from 8 to 32 and the tree depth from 1 to 9, we observe that the performance increases, and stabilizes for higher complexities. We provide figures of the data generated with these lower complexity hyperparameters in Appendix A.9 to give an impression. The correlation between performance and complexity supports our claim that learning complex decision boundaries is the driving force behind performance.
Figure 6: Ablation of the base size and maximum tree depth parameters of the Forest Dataset Generator. Figure shows normalized test accuracy of TabForest on the WhyTrees benchmark. A smaller model is used (dim: 512 -> 256, layers: 12 -> 8).
5.6 Superiority of Fine-tuning over Zero-shot
In the main results, we have seen that fine-tuning performs better than zero-shot. We look at this comparison in more detail. Figure 7(a) presents the performance of TabForestPFN on individual datasets from TabZilla. We can see clearly that fine-tuning strongly outperforms zero-shot when there are more than 10,000 observations. However, fine-tuning struggles on the two smallest datasets, which have four and six validation samples. Overall, fine-tuning outperforms zero-shot on 66% of the datasets. This percentage decreases to 54% for datasets smaller than a 1000 observations, and increases to 84% to datasets larger than a 1000 observations.
@@ 6 Conclusion
The introduction of TabPFN [20] has opened up a new field of in-context learning (ICL)-transformers for tabular data classification. Our research has demonstrated that these ICL-transformers learn to create complex decision boundaries. Furthermore, by fine-tuning and ensembling the pretraining dataset, we achieved performance levels that surpass XGBoost [6] and CatBoost [37] on third-party benchmarks [15, 32]. Although tree based-methods can continue to gain performance through hyperparameter optimization, we expect ICL-transformers can match their gains by scaling up the pretraining process.
Despite these advancements, there are still obstacles to overcome for ICL-transformers to fully replace tree-based methods in the realm of tabular data. One major challenge is the performance limitation of ICL-transformers due to GPU memory constraints. Our work uses fine-tuning as a solution to this problem, but it would be valuable to compare this approach to other concurrent research such as prompt tuning [11] and context distillation [30]. Moreover, our research solely focused on classification, although we expect a simple switch from cross-entropy loss to mean-squared-error loss would suffice to tackle regression tasks. Another area that requires exploration is the setting with an exceptionally high number of features [7], where the performance of ICL-transformers is unknown [31]. Lastly, tree-based methods have an edge in explaining the feature importance, and research needs to be done if ICL-transformers can achieve similar interpretability [40]. Overcoming these challenges will cement the ICL-transformer as the clear successor of tree-based methods.
@@ A.1 Ethics and Social Impact
A.2 Code Availability
Code is available at https://github.com/FelixdenBreejen/TabForestPFN
The code includes everything: downloading datasets, preprocessing result benchmarks, training the ICL-transformers, pretrained weights on Google Drive, notebooks with analysis, and an easy example to get you started with applying the TabForestPFN to your dataset.
The code is built upon the works of Grinsztajn et al., Hollmann et al. and McElfresh et al.. Although this resulted in a Frankenstein monster of a codebase, we have made great efforts to rewrite most of their code to integrate well. The code assumes a single server with access to 1 or more GPUs, with DistributedDataParallel used during pretraining and multiprocessing over different GPUs used during inference.
As currently there is no good tabular code base out there, we recommend anyone that is interested in doing research in tabular data to take a look at ours.
A.3 Data Preprocessing
@@
Because we fixed the input size of the neural network to $d_{f}=100$ features, we first select a subset of a $d_{f}$ features using scikit-learn’s SelectKBest [36]. If there are less than $d_{f}$ features, we add zeros to ensure exactly $d_{f}$ features. Following TabPFN, we scale by multiplying by $100/d_{f*}$, where $d_{f*}$ is the number of features after selecting a subset. To be robust to skewness and outliers, we transform the data using scikit-learn’s QuantileTransformer to follow a normal distribution. We make no distinction between numerical and categorical values in all our preprocessing.
@@
A.4 Training Settings
All ICL-transformer architectures, including the original TabPFN, use the same model. The model consists of 12 layers, 4 attention heads, a hidden dimension of 512, and 10 classes as output dimension. Pretraining uses batch size 64, learning rate 1e-4, weight decay 0.00, AdamW optimizer with betas (0.9, 0.95), cosine scheduling, and maximum global gradient norm 1.0. Fine-tuning is performed under batch size 1, learning rate 1e-5, weight decay 0.00, no scheduling, with early stopping, and a maximum of 300 steps. To fit the model on an RTX 3090 with 24GB during fine-tuning, we set the maximum support size to 8192 samples and the maximum query size to 1024. Pre-training uses data generated with maximum support size 1024 and maximum query size 128. We choose these settings because the maximum support size affects performance, but the maximum query size only affects inference speed.
On these settings, we train TabPFN and TabForest for 300,000 steps, while TabForestPFN trains for 600,000 steps. Both TabPFN and TabForest can likely train longer; however, we experienced high instability in performance after 300,000 steps depending on the seed. Sometimes, the fine-tuning ability of the model would completely collapse, and sometimes the performance would slowly overfit. We also chose 300,000 steps because we did not see improvement in fine-tuning performance in the steps afterward. We chose to use 600,000 steps for the combined TabForestPFN, because we believed double the data would allow us to use double the steps, and we experienced no instability on this setting. Running 600,000 steps on one RTX 3090 takes 24 GPU-days.
Choosing 300,000 steps does have a downside. In the results in section 5.2, we saw the retrained TabPFN and the TabForest have poor zero-shot performance. The number of steps therefore is a trade-off: a small number of steps is stable and has high fine-tuning performance, while a large number of steps has good zero-shot performance. Because in this paper we are interested in maximum performance, we prioritize the fine-tuned version on 300,000 steps.
A.5 Benchmark Metadata
Of both the TabZilla and the WhyTrees benchmark, we show the OpenML[47] datasets we use as well as their characteristics. See Table LABEL:table:metadata_tabzilla and Table LABEL:table:metadata_whytrees. The TabZilla table presents the 94 datasets picked out of the 176 total datasets.
@@ A.5 Benchmark Metadata
OpenML Observations Features Splits Classes
ID Name All Train Valid Test
44089 credit 16714 10000 2014 4700 10 2 2
44120 electricity 38474 10000 8542 19932 7 1 2
44121 covertype 566602 10000 50000 50000 10 1 2
44122 pol 10082 7057 907 2118 26 3 2
44123 house_16H 13488 9441 1214 2833 16 3 2
44125 MagicTelescope 13376 9363 1203 2810 10 3 2
44126 bank-marketing 10578 7404 952 $2$ 7 3 2
44128 MiniBooNE 72998 10000 18899 44099 50 1 2
44129 Higgs 940160 10000 50000 50000 24 1 2
44130 eye_movements 7608 5325 684 1599 20 3 2
44156 electricity 38474 10000 8542 19932 8 1 2
44157 eye_movements 7608 5325 684 1599 23 3 2
44159 covertype 423680 10000 50000 50000 54 1 2
45019 Bioresponse 3434 2403 309 722 419 5 2
45020 default-of-cred… 13272 9290 1194 2788 20 3 2
45021 jannis 57580 10000 14274 33306 54 1 2
45022 Diabetes130US 71090 10000 18327 42763 7 1 2
45026 heloc 10000 7000 900 2100 22 3 2
45028 california 20634 10000 3190 7444 8 1 2
45035 albert 58252 10000 14475 33777 31 1 2
45036 default-of-cred… 13272 9290 1194 2788 21 3 2
45038 road-safety 111762 10000 30528 50000 32 1 2
45039 compas-two-year… 4966 3476 447 1043 11 3 2

### v2

@@ 5.1 Introduction of the Benchmark Datasets
The TabZilla (McElfresh et al., 2023) benchmark tests 20 algorithms on 176 classification datasets with sizes ranging from 32 observations to over a million. We selected 94 out of 176 datasets, see appendix A.4. The medium-sized benchmark which we refer to as WhyTrees (Grinsztajn et al., 2022) consists of 23 classification datasets with 2923 to a maximum of 10,000 observations. The benchmark is split into 7 datasets with only numerical features and 16 datasets with both numerical and categorical features.
5.2 Main Results of TabForestPFN
@@
Zero-shot
TabScratch 0.000 0.000
TabPFN (original) 0.534 0.624
TabPFN (retrained) 0.481 0.635
TabForest 0.388 0.536
TabForestPFN 0.473 0.655
Fine-tuned
TabScratch 0.481 0.570
TabPFN (original) 0.742 0.775
TabPFN (retrained) 0.775 0.817
TabForest 0.853 0.854
TabForestPFN 0.842 0.849
The results on the TabZilla benchmark are shown in Table 3, see Appendix A.6 for alternative presentations. For the WhyTrees benchmark, the comparison of fine-tuned TabForestPFN with the benchmark algorithms is shown in Figure 4, and the comparison with other ICL-transformer variants in Table 2, while results on individual datasets can be seen in Appendix A.7. We present the running time of TabForestPFN on all datasets in Appendix A.5.
@@
TabForestPFN - Fine-tuned 1 27 8.4 7.0 0.840 0.902
TabPFN (retrained) - Fine-tuned 1 26 8.4 7.5 0.843 0.891
CatBoost 1 23 9.6 9.0 0.842 0.874
TabPFN (original) - Fine-tuned 1 26 9.6 10.0 0.834 0.902
TabForestPFN - Zero-shot 1 25 9.7 9.2 0.819 0.883
XGBoost 1 23 9.8 9.8 0.836 0.899
TabForest - Fine-tuned 1 27 10.9 10.0 0.806 0.873
TabPFN (retrained) - Zero-shot 1 26 11.2 10.5 0.797 0.858
LightGBM 1 27 11.8 12.2 0.787 0.867
TabPFN (original) - Zero-shot 1 26 12.1 12.0 0.777 0.841
RandomForest 1 26 12.1 12.0 0.792 0.851
Resnet 1 27 12.8 12.0 0.727 0.837
NODE 1 27 12.9 13.5 0.751 0.833
SAINT 1 27 13.1 13.8 0.729 0.803
SVM 1 26 13.3 14.0 0.710 0.801
FT-Transformer 1 24 13.6 13.2 0.733 0.805
TabScratch - Fine-tuned 1 25 13.6 13.0 0.752 0.819
DANet 2 27 15.6 16.0 0.718 0.769
TabForest - Zero-shot 3 26 15.7 16.0 0.709 0.822
MLP-rtdl 1 27 16.9 19.0 0.619 0.736
STG 1 27 17.1 19.0 0.592 0.664
LinearRegression 1 27 18.5 21.0 0.564 0.593
MLP 2 27 18.8 21.0 0.570 0.586
TabNet 2 27 19.1 20.2 0.579 0.666
DecisionTree 1 27 19.7 21.5 0.502 0.551
KNN 2 27 20.5 23.0 0.473 0.478
VIME 3 27 22.9 25.0 0.343 0.241
TabScratch - Zero-shot 28 28 28.0 28.0 0.000 0.000
@@
In these figures and tables, we see that fine-tuning considerably improves the performance of ICL-transformers. Where the zero-shot variants perform very mediocre compared to XGBoost and the other baselines, the fine-tuned variants are extremely competitive. Given the poor performance of TabScratch, we can clearly see that both the pretraining and the fine-tuning are important.
When comparing the pretraining datasets, we see that the best method differs by benchmark. Fine-tuned TabForest is the best on WhyTrees, while fine-tuned TabPFN is favored on TabZilla. As TabForest strength comes from generating complex decision boundaries, we conjecture that TabZilla has many datasets for which this property is not helpful. The zero-shot variants on both benchmarks clearly favor TabPFN, which is unsurprising given the unrealistic nature of the forest dataset generator.
The TabForestPFN combines the best of both worlds. For WhyTrees, we can see that the fine-tuned TabForestPFN has almost the same performance as TabForest, and for TabZilla, it has almost the same performance as TabPFN. We can also see that mixing in the forest dataset does not deteriorate the zero-shot performance on either benchmark, which establishes the TabForestPFN as the clear best method among the ICL-transformer variants.
@@ 5.3 Complexity of ICL-Transformers’ Decision Boundaries
In the previous section, we have seen that TabForest and TabPFN both achieve excellent performance as neural networks. Now we take a look at their decision boundaries. Repeating the analysis of the WhyTrees’ authors (Grinsztajn et al., 2022), we use the Electricity dataset (OpenML ID: 44156) to predict a binary target on two features.
To capture the complexity of the decision boundary, we define the complexity score $V$. We split the feature space into a total of $n$ grid cells where each cell has a predicted probability $p_{ij}$ for grid cell indices $(i,j)$. The complexity score $V$ is defined as the sum of absolute values between neighbor cells:
@@
We plot the results in Figure 1. We see that when fine-tuning, both TabPFN and TabForest can create decision boundaries that are more complex than their zero-shot variants. The complexity of the decision boundaries was one of the characteristics that explained why tree-based methods outperformed neural networks (Grinsztajn et al., 2022). These results suggest ICL-transformers can also create complexity in their decision boundaries.
In our intuition, the ICL-transformer learns how to create these decision boundaries during pretraining. We can interpret this from a weight initialization perspective. The weights of the pretrained ICL-transformer provide a good initialization for the model to create complex decision boundaries, while an ICL-transformer trained from scratch lacks this ability. For this reason, TabForest can create decision boundaries of higher complexity than TabPFN.
@@ 5.4 Ablation of the Forest Dataset Generator
In Section 4.1, we discussed two ways to influence the complexity of the forest dataset generator: the tree depth and the base size, which is the number of observations to fit the tree algorithm. We expect the performance of TabForest to increase when the complexity of the forest dataset generator increases.
In Figure 5 we show the results of pretraining different settings of base size and maximum tree depth on the WhyTrees benchmark. The tree depth is set to 1-25 as the base size changes, and the base size is fixed to 1024 as the tree depth changes. When scaling up the base size from 2 to 32 and the tree depth from 1 to 9, we observe that the performance increases, and stabilizes for higher complexities. We provide figures of the data generated with these lower complexity hyperparameters in Appendix A.9 to give an impression. The correlation between performance and complexity supports our claim that learning complex decision boundaries is the driving force behind the performance of fine-tuned TabForest.
Figure 5: Ablation of the base size and maximum tree depth parameters of the Forest Dataset Generator. Figure shows normalized test accuracy of TabForest on the WhyTrees benchmark.
5.5 Case Study of the Gap between Neural Networks and Tree Algorithms.
In Figure 6 we show the performance of fine-tuned TabPFN and TabForest on two specific datasets from WhyTrees. We selected these two datasets for the large gap between the neural networks (MLP, Resnet, SAINT, FT-Transformer) and the tree-based algorithms (XGBoost, GradientBoostingTree, RandomForest). The figure illustrates that ICL-transformers behave differently than other neural networks: their performance is closer to that of tree-based methods.
We propose the following explanation: these two datasets need highly complex decision boundaries. As tree-based methods are capable of creating these complex decision boundaries while neural networks struggle (Shah et al., 2020; Grinsztajn et al., 2022), fine-tuned ICL-transformers can also create them, as seen in Section 5.3. Furthermore, TabForest is naturally better at creating these decision boundaries than TabPFN, given that the forest dataset generator was specifically designed for this purpose. Figure 6 shows that TabForest significantly outperforms TabPFN on these two datasets. The explanation of the gap is further supported by Figure 1, as it illustrates the complexity of the decision boundaries for two variables from the Electricity dataset.
Figure 6: Comparison of fine-tuned TabForest and TabPFN on two datasets of the WhyTrees benchmark. These datasets show a big gap between neural networks and tree-based methods.
5.6 Improvement of Fine-tuning over Zero-shot
In the main results, we have seen that fine-tuning performs better than zero-shot. We look at this comparison in more detail. Figure 7(a) presents the performance of TabForestPFN on individual datasets from TabZilla. We can see clearly that fine-tuning strongly outperforms zero-shot when there are more than 10,000 observations. Overall, fine-tuning outperforms zero-shot on 57% of the datasets. This percentage decreases to 47% for datasets smaller than a 1000 observations and increases to 73% for datasets larger than a 1000 observations.
@@ 6 Conclusion
The introduction of TabPFN (Hollmann et al., 2023) has opened up a new field of in-context learning (ICL)-transformers for tabular data classification. Our research has demonstrated that fine-tuned ICL-transformers achieve excellent performance and also learn to create complex decision boundaries. Furthermore, by adding the forest dataset to the pretraining mixture, we achieved performance levels competitive with tree-based methods.
Despite these advancements, there are still obstacles for ICL-transformers to overcome if we want them to replace tree-based methods in the realm of tabular data. One major challenge is the performance limitation of ICL-transformers due to GPU memory constraints. Our work uses fine-tuning as a solution to this problem, but it would be valuable to compare this approach to other concurrent research such as prompt tuning (Feuer et al., 2024), in-context distillation (Ma et al., 2023) and retrieval (Thomas et al., 2024). Moreover, our research focused solely on classification, although we expect that a simple switch from cross-entropy loss to mean-squared-error loss would suffice to tackle regression tasks. Another area that requires exploration is the setting with an exceptionally high number of features (Cherepanova et al., 2024), where the performance of ICL-transformers is unknown (McCarter, 2024). Lastly, tree-based methods can explain which features are important for their predictions, and research is needed to determine if ICL transformers can achieve a similar feat (Rundel et al., 2024). Overcoming these challenges will cement the ICL-transformer as the clear successor to tree-based methods.
@@ A.1 Ethics and Social Impact
A.2 Data Preprocessing
@@
Because we fixed the input size of the neural network to $d_{f}=100$ features, we first select a subset of a $d_{f}$ features using scikit-learn’s SelectKBest Pedregosa et al. (2011). If there are less than $d_{f}$ features, we add zeros to ensure exactly $d_{f}$ features. Following TabPFN, we scale by multiplying by $100/d_{f*}$, where $d_{f*}$ is the number of features after selecting a subset. To be robust to skewness and outliers, we transform the data using scikit-learn’s QuantileTransformer to follow a normal distribution. We make no distinction between numerical and categorical values in all our preprocessing.
@@
A.3 Training Settings
All ICL-transformer architectures, including the original TabPFN, use the same model. The model consists of 12 layers, 4 attention heads, a hidden dimension of 512, and 10 classes as output dimension. Pretraining uses batch size 512, learning rate 1e-4, weight decay 0.00, AdamW optimizer with betas (0.9, 0.95), cosine scheduling, and maximum global gradient norm 1.0. Fine-tuning is performed under batch size 1, learning rate 1e-5, weight decay 0.00, no scheduling, with early stopping, and a maximum of 300 steps. To fit the model on an RTX 3090 with 24GB during fine-tuning, we set the maximum support size to 8192 samples and the maximum query size to 1024. Pre-training uses data generated with maximum support size 1024 and maximum query size 128. We choose these settings because the maximum support size affects performance, but the maximum query size only affects inference speed. On these settings, we train all ICL-transformers for 50,000 steps, which takes 4 GPU-days on one H100. Running all benchmarks takes one additional H100 GPU-day.
A.4 Benchmark Metadata
Of both the TabZilla and the WhyTrees benchmark, we show the OpenMLVanschoren et al. (2014) datasets we use as well as their characteristics. See Table LABEL:table:metadata_whytrees and Table LABEL:table:metadata_tabzilla. The TabZilla table presents the 94 datasets picked out of the 176 total datasets.
@@ A.4 Benchmark Metadata
OpenML Observations Features Splits Classes
ID Name All Train Valid Test
44089 credit 16714 10000 2014 4700 10 2 2
44120 electricity 38474 10000 8542 19932 7 1 2
44121 covertype 566602 10000 50000 50000 10 1 2
44122 pol 10082 7057 907 2118 26 3 2
44123 house_16H 13488 9441 1214 2833 16 3 2
44125 MagicTelescope 13376 9363 1203 2810 10 3 2
44126 bank-marketing 10578 7404 952 $2$ 7 3 2
44128 MiniBooNE 72998 10000 18899 44099 50 1 2
44129 Higgs 940160 10000 50000 50000 24 1 2
44130 eye_movements 7608 5325 684 1599 20 3 2
44156 electricity 38474 10000 8542 19932 8 1 2
44157 eye_movements 7608 5325 684 1599 23 3 2
44159 covertype 423680 10000 50000 50000 54 1 2
45019 Bioresponse 3434 2403 309 722 419 5 2
45020 default-of-cred… 13272 9290 1194 2788 20 3 2
45021 jannis 57580 10000 14274 33306 54 1 2
45022 Diabetes130US 71090 10000 18327 42763 7 1 2
45026 heloc 10000 7000 900 2100 22 3 2
45028 california 20634 10000 3190 7444 8 1 2
45035 albert 58252 10000 14475 33777 31 1 2
45036 default-of-cred… 13272 9290 1194 2788 21 3 2
45038 road-safety 111762 10000 30528 50000 32 1 2
45039 compas-two-year… 4966 3476 447 1043 11 3 2
@@
A.5 Run times
Table LABEL:table:runtime_whytrees and LABEL:table:runtime_tabzilla present the run times of TabForestPFN on both the WhyTrees and the TabZilla benchmark. All fine-tuning runs take at most 220 seconds per cross validation split, with an average of 68 seconds. Runtimes differ by GPU, and creating a fair comparison with CPU-based methods is difficult due to the different hardware used. As main takeaway, we recommend to summarize the fine-tuning run time as ”a few minutes at most for a dataset of around 10,000 observations”.
Table 6: Run times of TabForestPFN of the WhyTrees Benchmark. The runtime is the end-to-end time in seconds for one cross validation split. End-to-end time includes loading, preprocessing, training and testing.
OpenML Size Run time (s)
ID Name Obs. Feat. Zero-shot Fine-tuned
44089 credit 10000 10 9 103
44120 electricity 10000 7 15 151
44121 covertype 10000 10 34 167
44122 pol 7057 26 6 57
44123 house_16H 9441 16 8 72
44125 MagicTelescope 9363 10 7 105
44126 bank-marketing 7404 7 7 68
44128 MiniBooNE 10000 50 28 126
44129 Higgs 10000 24 34 119
44130 eye_movements 5325 20 5 63
44156 electricity 10000 8 17 142
44157 eye_movements 5325 23 6 65
44159 covertype 10000 54 37 219
45019 Bioresponse 2403 419 8 34
45020 default-of-credit-card-clients 9290 20 7 81
45021 jannis 10000 54 23 130
45022 Diabetes130US 10000 7 25 95
45026 heloc 7000 22 6 56
45028 california 10000 8 11 112
45035 albert 10000 31 21 103
45036 default-of-credit-card-clients 9290 21 8 79
45038 road-safety 10000 32 30 153
45039 compas-two-years 3476 11 5 43
Table 7: Run times of TabForestPFN of the TabZilla Benchmark. The runtime is the end-to-end time in seconds for one cross validation split. End-to-end time includes loading, preprocessing, training and testing.
OpenML Size Run time (s)
ID Name Obs. Feat. Zero-shot Fine-tuned
3 kr-vs-kp 2556 36 4 29
4 labor 45 16 3 13
9 autos 163 25 3 11
10 lymph 118 18 3 9
11 balance-scale 499 4 3 32
12 mfeat-factors 1600 216 5 26
14 mfeat-fourier 1600 76 4 29
15 breast-w 559 9 3 19
16 mfeat-karhunen 1600 64 4 22
18 mfeat-morphological 1600 6 4 22
23 cmc 1177 9 4 20
25 colic 294 26 3 10
27 colic 294 22 3 11
29 credit-approval 552 15 4 22
30 page-blocks 4377 10 5 40
35 dermatology 292 34 3 13
37 diabetes 614 8 3 19
39 sonar 166 60 3 11
40 glass 170 9 3 10
43 spambase 3680 57 6 42
45 splice 2552 60 3 33
47 tae 120 5 3 11
48 heart-c 241 13 3 11
49 tic-tac-toe 766 9 3 20
50 heart-h 234 13 2 12
53 vehicle 676 18 3 23
59 iris 120 4 3 16
2074 satimage 5144 36 6 55
2079 eucalyptus 588 19 3 18
2867 anneal 718 38 3 26
3485 scene 1925 299 6 37
3512 synthetic_control 480 60 3 18
3540 analcatdata_boxing1 96 3 3 12
3543 irish 400 5 4 19
3549 analcatdata_authorship 672 70 4 25
3560 analcatdata_dmft 637 4 3 20
3561 profb 536 9 3 16
3602 visualizing_environmental 88 3 3 10
3620 fri_c0_100_5 80 5 3 13
3647 rabe_266 96 2 3 14
3711 elevators 13279 18 9 101
3731 visualizing_livestock 104 2 3 15
3739 analcatdata_chlamydia 80 3 3 16
3748 transplant 104 3 3 12
3779 fri_c3_100_5 80 5 3 13
3797 socmob 924 5 3 19
3896 ada_agnostic 3648 48 6 36
3902 pc4 1166 37 4 23
3903 pc3 1249 37 4 26
3904 jm1 8707 21 6 117
3913 kc2 416 21 3 26
3917 kc1 1687 21 4 48
3918 pc1 887 21 3 16
3953 adult-census 26048 14 14 175
9946 wdbc 455 30 4 17
9952 phoneme 4322 5 4 44
9957 qsar-biodeg 843 41 4 23
9960 wall-robot-navigation 4364 24 5 42
9964 semeion 1273 256 5 26
9971 ilpd 465 10 4 20
9978 ozone-level-8hr 2026 72 4 25
9984 fertility 80 9 3 13
10089 acute-inflammations 96 6 3 10
10093 banknote-authentication 1096 4 4 20
10101 blood-transfusion-service-center 598 4 3 20
14952 PhishingWebsites 8843 30 8 103
14954 cylinder-bands 432 37 4 16
14965 bank-marketing 36168 16 17 165
14967 cjs 2236 33 4 79
125920 dresses-sales 400 12 4 18
125921 LED-display-domain-7digit 400 7 4 16
145793 yeast 1015 8 4 19
145799 breast-cancer 228 9 3 11
145836 blood-transfusion-service-center 598 4 3 21
145847 hill-valley 968 100 4 47
145977 ecoli 268 7 3 12
145984 ionosphere 280 34 3 12
146024 lung-cancer 24 56 3 14
146063 hayes-roth 128 4 3 14
146065 monks-problems-2 480 6 2 22
146192 car-evaluation 1382 21 4 27
146210 postoperative-patient-data 70 8 3 13
146607 SpeedDating 6702 120 6 57
146800 MiceProtein 864 77 4 28
146817 steel-plates-fault 1552 27 4 22
146818 Australian 552 14 4 23
146820 wilt 3871 5 4 30
146821 car 1382 6 4 30
167140 dna 2548 180 4 26
167141 churn 4000 20 5 41
167211 Satellite 4080 36 5 40
168911 jasmine 2386 144 4 36
190408 Click_prediction_small 31958 11 14 129
360948 libras 288 104 3 11